Andrew and Joseph took a trip to Vintage Vinyl Record Store at the store Andrew bought 6 records and 3 CDs for a total of $85.23. Joseph bought 4 records and 3 CDs for a total of $67.81


please help me write and solve a system of linear equations to find the cost of an individual record and an individual CD?
thank you!

Answer by stanbon(75887)

Andrew and Joseph took a trip to Vintage Vinyl Record Store at the store Andrew bought 6 records and 3 CDs for a total of $85.23. Joseph bought 4 records and 3 CDs for a total of $67.81
--------------
6r + 3c = 85.23
4r + 3c = 67.81
------------------------
Subtract the 2nd equation from the 1st and solve for "r":
2r = 17.42
r = $8.71 (price of one record)
-----
Substitute into one of the original equations and solve for "c":
4(8.71) + 3c = 67.81
c = $10.99 (price of one CD)
===================================
Cheers,
Stan H.